# Enumerate–Conjecture–Prove: Formally Solving Answer-Construction Problems in Math Competitions

### v3

@@ B.1 Prompt for Formalization
‘‘‘lean
your code
B.2 Prompt for Enumerator Coding
LLM Prompt I have a difficult high-school competition-level math problem. Your task is to write a Python program to enumerate possible answers to assist in conjecturing and proving the true answer. Use the following format for the delimited output:
‘‘‘python
your code
B.3 Prompt for Conjecturing
LLM Prompt I have a difficult high-school competition-level math problem in Lean, which asks to find a answer set with a mathematical proof. Your task is to clearly state the final answer in Lean exprssion with explanations.You do not have to derive the whole proof. I have enumerated the answer(s) via program as a hint for you, and you may look at the pattern in it to help propose the final answer. Your final answer should be in closed form in Lean and generalized from enumerated answer that can eliminate the existential quantifier. You must not echo the problem statement in your answer. Use the following format for the delimited output:
‘‘‘lean
your code

### v4

@@ Appendix C Human Evaluation Examples
Analysis. The logical shape is completely wrong. The problem asks for all possible values of $a+b+c+d$ that occur for some positive integers satisfying the equation. However, the formalized theorem states that for some specific $a,b,c,d>0$ the equation holds iff $a+b+c+d\in answer$, which would mean the equation is true for those exact $a,b,c,d$ precisely when the sum is not square-free.
Appendix D Full Tables
D.1 ConstructiveBench Full (3640 problems)
DeepSeek-Prover-V2-7B Goedel-Prover-SFT Kimina-Prover-Preview-7B
Conjecturer CoT ECP Union CoT ECP Union CoT ECP Union
GPT-5 nano 28.8% 30.9% 34.5% 24.9% 26.1% 29.5% 22.3% 23.2% 27.4%
GPT-5 mini 32.5% 33.1% 35.0% 27.2% 27.6% 29.1% 25.0% 25.2% 27.6%
DeepSeek-V3 10.1% 18.0% 18.7% 7.0% 15.0% 15.3% 7.8% 14.0% 14.8%
DeepSeek-V3.1 31.1% 32.7% 34.3% 26.2% 27.5% 29.0% 23.8% 25.1% 27.6%
GPT-4.1 mini 7.3% 17.0% 17.7% 5.3% 11.1% 12.0% 6.0% 13.1% 13.9%
GPT-4.1 nano 9.8% 14.3% 16.7% 7.1% 10.0% 11.6% 7.9% 11.2% 13.1%
Table 4: ConstructiveBench full end-to-end results (Pass@32)
D.2 ConstructiveBench After-Cutoff (106 problems)
DeepSeek-Prover-V2-7B Goedel-Prover-SFT Kimina-Prover-Preview-7B
Conjecturer CoT ECP Union CoT ECP Union CoT ECP Union
GPT-5 nano 13.2% 15.1% 18.9% 11.3% 12.3% 14.2% 10.4% 11.3% 13.2%
GPT-5 mini 17.9% 22.6% 22.6% 15.1% 16.0% 17.9% 13.2% 14.2% 15.1%
DeepSeek-V3 2.8% 5.7% 5.7% 1.9% 4.7% 4.7% 1.9% 3.8% 3.8%
DeepSeek-V3.1 17.0% 17.9% 20.8% 15.1% 16.0% 17.9% 13.2% 14.2% 15.1%
GPT-4.1 mini 4.7% 6.6% 6.6% 3.8% 5.7% 5.7% 3.8% 4.7% 4.7%
GPT-4.1 nano 2.8% 4.7% 5.7% 1.9% 2.8% 3.8% 1.9% 2.8% 3.8%
Table 5: ConstructiveBench After-Cutoff end-to-end results (Pass@32)
D.3 PutnamBench Subset (337 problems)
DeepSeek-Prover-V2-7B Goedel-Prover-SFT Kimina-Prover-Preview-7B
Conjecturer CoT ECP Union CoT ECP Union CoT ECP Union
GPT-5 nano 1.2% 1.5% 1.8% 0.9% 0.9% 1.2% 0.9% 1.2% 1.5%
GPT-5 mini 1.2% 1.8% 1.8% 0.9% 0.3% 0.9% 0.9% 0.6% 0.9%
DeepSeek-V3 0.3% 0.9% 0.9% 0.0% 0.3% 0.3% 0.0% 0.0% 0.0%
DeepSeek-V3.1 1.2% 1.2% 1.5% 0.6% 0.3% 0.6% 0.6% 0.6% 0.9%
GPT-4.1 mini 0.6% 0.6% 0.9% 0.6% 0.6% 0.9% 0.0% 0.0% 0.0%
GPT-4.1 nano 0.9% 0.6% 0.9% 0.3% 0.0% 0.3% 0.0% 0.0% 0.0%
Table 6: PutnamBench subset end-to-end results (Pass@32) with timeout=120s.
D.4 Full Answer-Construction Accuracies
Conjecturer ConstructiveBench (Full) ConstructiveBench (After-Cutoff) PutnamBench
CoT / ECP Union CoT / ECP Union CoT / ECP Union
GPT-5 nano 45.1% / 62.3% 66.9% 28.3% / 55.7% 57.6% 16.3% / 32.3% 36.8%
GPT-5 mini 69.7% / 73.6% 78.9% 51.9% / 58.5% 64.2% 54.9% / 57.6% 71.8%
DeepSeek-V3 17.3% / 40.7% 43.4% 2.8% / 26.4% 26.4% 19.0% / 31.8% 36.8%
DeepSeek-V3.1 65.4% / 75.0% 79.2% 44.3% / 67.9% 69.8% 53.7% / 62.3% 73.6%
GPT-4.1 mini 16.6% / 36.5% 37.8% 11.3% / 23.6% 24.5% 15.1% / 23.2% 29.4%
GPT-4.1 nano 15.3% / 28.5% 33.0% 4.7% / 18.9% 20.8% 7.7% / 11.6% 17.2%
Table 7: Answer-construction accuracy. Note that, DeepSeek-V3.1 has knowledge cutoff after June 2024, so this model’s ConstructiveBench (After-Cutoff) accuracy may be contaminated.